40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5989     Entry time: Wed Nov 23 16:48:39 2011
Author: Suresh 
Type: Update 
Category: General 
Subject: cable cleanup 

[Koji Suresh]

As part of the general lab clean up we removed many unused BNC cables (long and short) from around the SP table.  We removed one very long BNC cable which was connected on one side to an PEM input and not connected on the other side near the 1X2 rack..   There were several cables from an old SURF phase camera project which were still attached to a couple of RF amps on the SP tables and running towards the 1X6 rack. 

We also removed some unused power cables  plugged into a power distribution strip near Megatron.

 

ELOG V3.1.3-